OpenStudy (anonymous):

Solve the following inequality and graph the solution set |x+6|

OpenStudy (anonymous):

|x+6|< or = 5 -5<|x + 6|<5 -6-5<lx|< 5 - 6 -11<x<-1 Therefore [-11, -1] is the answer \[\le\] This means less than or equal to hence why I used [ brackets instead of ( brackets.

OpenStudy (anonymous):

I didn't know how to make the line go under it on here I knew what it meant :)

OpenStudy (saifoo.khan):

you can use the "equation editor" here to use it.

OpenStudy (saifoo.khan):

\[\le,\ge\]

OpenStudy (anonymous):

O ok I see how to do it now thanks